OpenStudy (anonymous):

Use Green's Theorem to evaluate \[\int\limits_{C}^{}xydx+x ^{2}y ^{3}dy\] where C is the triangle with vertices (0,0), (1,0), and (1,2).

OpenStudy (anonymous):

Take your partial derivatives: \[N_x=\frac{\partial (x^2y^3)}{\partial x}=\cdots\\ M_y=\frac{\partial (xy)}{\partial y}=\cdots\] Then by Green's theorem, you have \[\int_C M~dx+N~dy=\int\int_D(N_x-M_y)~dA\] where \(D\) is the triangular region defined by \[D:=\left\{(x,y)~:~0\le x\le1,~0\le y\le2x\right\}\]